OpenStudy (anonymous):

The profit function p(x) of a tour operator is modeled by p(x) = −2x2 + 700x − 10000 where x is the average number of tours he arranges per day. What is the range of the average number of tours he must arrange per day to earn a monthly profit of at least $50,000? between 150 and 200, exclusive more than 200 between 150 and 200, inclusive less than 150

OpenStudy (loser66):

Any idea?

OpenStudy (anonymous):

set the equation equal to profit per day then solve the equation for x

OpenStudy (anonymous):

i have no idea how to do that.. p(100)=−2(1002)+700(100)−10000=−20000+70000−10000=40,000..? number 40000 equals f(100), which is an example of how to calculate f(x) for x=100.. that's all I know